OpenStudy (rock_mit182):

IDk how tostart please help! int \frac{ 8dx }{ (4x^2+1)^2 } int \csc(x) sin(3x) dx

OpenStudy (rock_mit182):

\[\int\limits \frac{ 8dx }{( 4x^2+1)^2 }\]

OpenStudy (rock_mit182):

\[\int\limits \csc(x) \sin(3x)dx\]

OpenStudy (amorfide):

for the first one, bring the bottom to the top to get a negative power then you can just use the trick of \[\int\limits_{}^{}2(2x+3)^3 dx=\] would be add one to the power divide by the power, then you also divide by the differential of the inside of the bracket to get \[\frac{ 2 }{ 4 \times 2 }(2x+3)^{4}\]

OpenStudy (amorfide):

use the same logic

OpenStudy (rock_mit182):

@phi

OpenStudy (rock_mit182):

is that easy are you sure ?

OpenStudy (amorfide):

yes

OpenStudy (rock_mit182):

what a relief

OpenStudy (amorfide):

or you could split into partial fractions

OpenStudy (rock_mit182):

im going to try both methos

OpenStudy (rock_mit182):

why not ?

OpenStudy (amorfide):

okay I will work on the second integral

OpenStudy (amorfide):

you can try the partial fraction method if you want

OpenStudy (amorfide):

the last one is integrating by parts

OpenStudy (rock_mit182):